# The NTU Partitioned Matching Game for International Kidney Exchange Programs

### v1

@@ 1 Introduction
For a positive integer $k$, we use the notation $[k]=\{1,2,\dots,k\}$. The NTU partitioned matching game is defined by a graph $G=(V;E)$ and a partition $V=V_{1}\cup V_{2}\cup\dots\cup V_{m}$ of $V$, where $m$ is the number of players. Each class of this partition corresponds to a player of the game; we say that the vertices in $V_{i}$ belong to player $i\in[m]$ (sometimes we identify the player and the set and say that player $i$ has size $|V_{i}|$). If each $V_{i}$ has size 2, then the problem is called the NTU matching game with couples.
@@
Given a matching $M$ in $G$, the utility of $M$ for player $i$ is defined by $u_{i}(M)=|V(M)\cap V_{i}|$, where $V(M)$ denotes the vertex set of $M$, i.e. the utility is the number of vertices of $V_{i}$ covered by $M$.
@@
A coalition $\{i_{1},i_{2},\dots,i_{k}\}$ of players is strongly blocking for the matching $M$ if there exists a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $u_{i_{j}}(M^{\prime})>u_{i_{j}}(M)$ for every $j\in[k]$. Similarly, a coalition $\{i_{1},i_{2},\dots,i_{k}\}$ of players is weakly blocking for a matching $M$ if there exists a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $u_{i_{j}}(M^{\prime})\geq u_{i_{j}}(M)$ for every $j\in[k]$, and $u_{i_{j}}(M^{\prime})>u_{i_{j}}(M)$ for at least one $j\in[k]$.
@@
The Gallai-Edmonds decomposition
The Gallai-Edmonds (GE) decomposition of a graph $G=(V,E)$ consists of three disjoint sets of vertices $A(G),C(G),D(G)$ with $V=A(G)\cup C(G)\cup D(G)$ and the following properties:
  • $C(G)$ is the set of even-sized components and $D(G)$ is the set of odd-sized components in the graph $G\setminus A(G)$,
  • Any maximum size matching covers all but $|D(G)|-|A(G)|$ vertices, which are all members of $D(G)$ and are in different components of $G[D(G)]$,
  • Every maximum size matching has the following structure: it induces a perfect matching in $G[C(G)]$, it induces an almost perfect matching in each odd component in $G[D(G)]$ (meaning at most one vertex is exposed) and every vertex $v\in A(G)$ is matched to some vertex in $D(G)$.
  • Every component $Q$ of $G[D(G)]$ is hypomatchable, that is, $Q-v$ has a perfect matching for every vertex $v$ of $Q$. Furthermore, there is a maximum size matching where $v$ is exposed.
Matroids
In Sections 2 and 3, we will use some basic facts from matroid theory. A matroid is a pair $(S,\mathcal{I})$ of a finite set $S$ and a nonempty family $\mathcal{I}\subseteq 2^{S}$ satisfying the axioms (i) $A\subseteq B\in\mathcal{I}$ implies $A\in\mathcal{I}$, and (ii) for any $A,B\in\mathcal{I}$ with $|A|<|B|$, there is an element $x\in B\setminus A$ with $A+x\in\mathcal{I}$. A set in $\mathcal{I}$ is called an independent set. By axiom (ii), all inclusion-wise maximal independent sets have the same size; these are called the bases, and their size is the rank of the matroid.
We consider two types of matroids. Given a partition $S_{1},\dots,S_{m}$ of a ground set $S$ and numbers $0\leq q_{i}\leq|S_{i}|$, the corresponding generalized partition matroid is defined as =$\mathcal{I}=\{X\subseteq S:|X\cap S_{i}|\leq q_{i}\ (i\in[m]\}$. The rank of this matroid is $\sum q_{i}$. We also define a class of matroids using matchings. Given a graph $G=(S,E)$, the matching matroid of $G$ is defined as =$\mathcal{I}=\{X\subseteq S:\exists\ \text{matching $M$ in $G$ such that }X\subseteq V(M)\}$. The rank of this matroid is twice the maximum size of a matching. See [15, Chapter 39] for further details. We use the fact that a maximum-size common independent set of two matroids can be found in polynomial time. This implies the following lemma that we will use later.
@@
Given a graph $G=(V;E)$, a partition $V=V_{1}\cup V_{2}\cup\dots\cup V_{m}$ of $V$, and a vector $q\in\mathbb{Z}_{+}^{m}$, it can be decided in polynomial time if there is a matching $M$ such that $|V(M)\cap V_{i}|\geq q_{i}$ for every $i\in[m]$.
@@
Let $\mathcal{M}_{1}$ be the generalized partition matroid given by a partition $V_{1},V_{2},\dots,V_{m}$ and numbers $q_{1},\dots,q_{m}$, and let $\mathcal{M}_{2}$ be the matching matroid of $G$. There is a common independent set of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ of size $\sum_{i=1}^{m}q_{i}$ if and only if there is a matching $M$ with $|V(M)\cap V_{i}|\geq q_{i}$ for every $i\in[m]$. ∎
@@ 2 The NTU matching game with couples
In this section, we study the case where every partition class has size at most 2. We show that in this case all four studied problems become efficiently solvable. However, deciding the emptiness of the strong core still requires a complicated approach relying on many structural results.
First of all, we may assume without loss of generality that every partition class has size exactly two; otherwise we add an isolated vertex and make it belong to that player. With this assumption, the partition classes can be considered as edges, which we call player edges. The perfect matching defined by the player edges is denoted by $M_{0}$. Sometimes we identify a player and the corresponding player edge, so the edge of player $A$ is also denoted by $A$. These edges are not part of the original graph $G=(V,E)$, although $E$ can also contain edges between the two vertices of a player.
We say that a path or cycle in $E\cup M_{0}$ is $M_{0}$-alternating if every other edge is in $M_{0}$. Note that if $uv$ is both in $E$ and in $M_{0}$, then these two edges form an $M_{0}$-alternating cycle. Given an $M_{0}$-alternating path or cycle $P$, we can consider the coalition formed by $P\cap M_{0}$, and the matching on the vertices of the coalition given by $P\cap E$. We say that $P$ is a (weakly/strongly) blocking path or cycle for a given matching $M$ if $P\cap M_{0}$ is a weakly/strongly blocking coalition for $M$ as witnessed by the matching $P\cap E$.
@@ 2.1 Weak core
As we shall see, the problems concerning the weak core are considerably simpler than those concerning the strong core. We begin with a characterization result.
@@
If $M$ is in the weak core, then conditions a) and b) hold, since otherwise the $M_{0}$-alternating path or cycle would block $M$. Suppose a) and b) hold, but $M$ is not in the weak core. There is a coalition of players with a matching $M^{\prime}$ such that every player in the coalition is strictly better off with $M^{\prime}$, therefore those players who have both of their vertices covered by $M$ cannot belong to the coalition – we can delete these players from the graph. Take the symmetric difference of the remaining player edges with $M^{\prime}$. This consists of alternating cycles and paths. If there is an alternating cycle in the symmetric difference, then it contradicts a). If there is an alternating path, then the player edges at the two ends of the path have one vertex covered by $M^{\prime}$, therefore they have 0 vertices covered by $M$, which contradicts b). ∎
@@
If every player has size 2, then we can decide in polynomial time if a given matching $M$ is in the weak core.
@@
We can check the conditions of Lemma 2 in polynomial time. We restrict the graph to the players who have at most one vertex covered by $M$.
We can check if a given player edge is in an alternating cycle, since it is equivalent to checking if, after the deletion of this edge, the remaining graph (the original graph $G$ together with the player edges) has a perfect matching. Hence, condition a) can be checked by iterating through the players.
We can also check whether there is an alternating path between two given player edges, knowing that these do not belong to an alternating cycle, since this is equivalent to checking if by deleting these two edges the remaining graph has a matching that covers all but two vertices (an almost perfect matching). We can do this for any two player edges that have 0 vertices covered by $M$. ∎
In fact, we can show that the weak core of the NTU matching game with couples is never empty.
@@
If every player has size 2, then the weak core is not empty.
@@
We can construct a matching $M^{*}$ in the weak core the following way. Let $M_{0}$ be the perfect matching formed by the player edges as before. We check if there is an $M_{0}$-alternating cycle in the current graph. If there is, then let the edges of $E$ in the cycle belong to $M$, and delete the vertices of the cycle from the graph. Repeat this with the remaining graph, until there are no $M_{0}$-alternating cycles. Let $M$ denote the resulting matching, and let $M^{*}$ be a maximum size matching in $G$ that covers every vertex of $M$. We claim that $M^{*}$ is in the weak core, since the conditions of Lemma 2 are met. Condition a) clearly holds, since the construction deleted players that were covered twice by $M$ (and hence by $M^{*}$), and the remaining graph did not contain an $M_{0}$-alternating cycle.
@@ 2.2 Strong core
In the NTU matching game with couples, a matching $M$ is in the strong core if and only if
@@
    every player that is in an $M_{0}$-alternating cycle has both its vertices covered by $M$,
@@
    there is no $M_{0}$-alternating path from a player that has none of its vertices covered by $M$ to a player that has at most one vertex covered by $M$,
@@
It is easy to see that if $M$ is in the strong core, then these conditions hold, since otherwise the $M_{0}$-alternating path or cycle would block $M$. Suppose that the conditions hold, but the matching $M$ is not in the strong core. Then there is a weakly blocking coalition of players, together with a matching $M^{\prime}$, such that there is one player that is better off with $M^{\prime}$ than $M$, and the others are not worse off. Take the symmetric difference of $M^{\prime}$ and $M_{0}$. This is a disjoint union of $M_{0}$-alternating cycles and paths. If the player that is better off by $M^{\prime}$ has 0 vertices covered by $M$, then it has at least one vertex covered by $M^{\prime}$, so it is either in an $M_{0}$-alternating cycle, which contradicts condition a), or it is in an $M_{0}$-alternating path. In the latter case, the two player edges at the ends of the path are players who have one vertex covered by $M^{\prime}$, so they have at most one vertex covered by $M$, which contradicts condition b). If the player that is better off by $M^{\prime}$ has 1 vertex covered by $M$, then it has two vertices covered by $M^{\prime}$, so it is either in an $M_{0}$-alternating cycle, which contradicts condition a), or it is in an $M_{0}$-alternating path (and it is not at the end of the path), but the two player edges at the ends of the path are players that have one vertex covered by $M^{\prime}$, so they have at most one vertex covered by $M$, which contradicts condition b) or c). ∎
@@
In the the NTU matching game with couples, we can decide in polynomial time if a given matching $M$ is in the strong core.
@@
We need to show that we can check the conditions given by Lemma 5 in polynomial time. We can check if a given player edge is in an $M_{0}$-alternating cycle, since it is equivalent to checking if, after the deletion of this edge, the remaining graph (the original graph $G$ together with the player edges) has a perfect matching. We can do this for every player edge with at most one vertex covered by $M$.
@@
We can check whether there is an $M_{0}$-alternating path between two given player edges, knowing that these do not belong to an $M_{0}$-alternating cycle, since this is equivalent to checking if by deleting these two edges the remaining graph has an almost perfect matching. We can do this for any two player edges such that one of them has 0 vertices covered by $M$, and the other one has at most one vertex covered by $M$.
@@
For three given player edges, such that we know that none of them belong to an $M_{0}$-alternating cycle, we can check if there is an alternating path that contains all three of them, because this is equivalent to checking if by deleting all three of these edges, the remaining graph has an almost perfect matching. We can check this for any three player edges that have one vertex covered by $M$. ∎
2.3 Existence of a strong core matching

### v2

@@ 1.5 Preliminaries
For a positive integer $k$, we use the notation $k]=\{1,2,\dots,k\}$. An instance $I$ of the NTU Partitioned Matching Game is defined by a tuple $(G,\Pi)$, where $G=(V;E)$ is a graph and $\Pi=(V_{1},V_{2},\dots,V_{m})$ is a partition of $V$ into $m$ disjoint subsets, where $m$ is the number of players. Each class of this partition corresponds to a player; we say that the vertices in $V_{i}$ belong to player $i\in[m]$ (sometimes we identify the player and the set and say that player $i$ has size $|V_{i}|$). If each $V_{i}$ has size 2, then the problem is called the NTU matching game with couples.
@@
Given a matching $M$ in $G$, the utility of $M$ for player $i$ is defined by $u_{i}(M)=|V(M)\cap V_{i}|$, where $V(M)$ denotes the vertex set of $M$; i.e., the utility is the number of vertices of $V_{i}$ covered by $M$.
@@
A coalition $\mathcal{P}=\{i_{1},i_{2},\dots,i_{k}\}$ of players is strongly blocking for the matching $M$ if there exists a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $u_{i_{j}}(M^{\prime})>u_{i_{j}}(M)$ for every $j\in[k]$. Similarly, a coalition $\mathcal{P}=\{i_{1},i_{2},\dots,i_{k}\}$ of players is weakly blocking for a matching $M$ if there exists a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $u_{i_{j}}(M^{\prime})\geq u_{i_{j}}(M)$ for every $j\in[k]$, and $u_{i_{j}}(M^{\prime})>u_{i_{j}}(M)$ for at least one $j\in[k]$.
@@
Given an instance $I=(G,\Pi)$ of the NTU partitioned matching game with partition classes of size at most $\ell$, we denote the problems of deciding the existence of a weak core or strong core matching by NTU-PMG-WC-Exist-$\ell$ and NTU-PMG-SC-Exist-$\ell$, respectively. Similarly, we denote the problem of deciding if a given matching $M$ in $I$ is in the weak or strong core by NTU-PMG-WC-Verif-$\ell$ and NTU-PMG-SC-Verif-$\ell$, respectively. The $\ell=2$ case corresponds to the NTU matching game with couples.
The following lemma is a central tool for finding matchings in the weak or strong cores, and for verifying if a given matching is in the weak or strong core.
@@
Given a graph $G=(V;E)$, a partition $U_{1}\cup U_{2}\cup\dots\cup U_{l}$ of $V$, and a vector $q\in\mathbb{Z}_{+}^{m}$ with $q_{i}\leq|U_{i}|$, it can be decided in $\mathcal{O}(|V|^{2.5})$ time if there is a matching $M$ such that $|U_{i}\cap V(M)|\geq q_{i}$ for every $i\in[l]$. Furthermore, given a cost function $c:E\to\mathbb{R}$, we can find a minimum cost such matching in $\mathcal{O}(|V|^{4})$ time, if one exists.
@@
[…]
To find a minimum cost matching $M$ with $|V(M)\cap U_{i}|\geq q_{i}$ $\forall i\in[l]$, if one exists, given a cost function $c:E\to\mathbb{R}$, let $R:=\max\{|c(e)|\mid e\in E\}+1$. It suffices to find a minimum cost perfect matching in $\hat{G}$ with the cost function $c^{\prime}(e):=c(e)$, if $e\in E$ and $c^{\prime}(e)=0$ for $e\in\hat{E}\setminus E$.
[…]
@@ 2 The NTU matching game with couples
In this section, we study the case where every partition class has size at most 2. We show that in this case, deciding membership and deciding existence are efficiently solvable for both the weak and strong cores, although deciding the emptiness of the strong core requires a complicated approach relying on many structural results. Somewhat surprisingly, the computational difficulty of optimization is the other way around: it is NP-hard for the weak core, while polynomial-time solvable for the strong core.
The set $M_{0}$ of player edges.
First of all, we may assume without loss of generality that every partition class has size exactly two; otherwise we add an isolated vertex and make it belong to that player. With this assumption, the partition classes can be considered as graph edges, which we call player edges. The perfect matching on vertex set $V$ defined by the player edges is denoted by $M_{0}$. For a player $A$, we will denote the corresponding player edge by $e_{A}$. These edges are not part of the original graph $G=(V,E)$, although $E$ can also contain an edge $e=aa^{\prime}$ between the two vertices of a player $A=\{a,a^{\prime}\}$, in which case $e_{A}$ will be a parallel edge to $e$. See Figure 1 for an illustration.
Figure 1: An example of the graph supplemented with the player edges. The dashed edges denote $E$ and the solid edges denote the player edges $M_{0}$.
We say that a path or cycle in $E\cup M_{0}$ is $M_{0}$-alternating if it alternates between edges in $M_{0}$ and edges in $E$. Note that if both $E$ and $M_{0}$ contain edges between vertices $u$ and $v$, then these two parallel edges form an $M_{0}$-alternating cycle.
For an $M_{0}$-alternating path or cycle $P$, we can consider the coalition formed by $\mathcal{P}=\{i\in[m]\mid e_{i}\in P\}$, and the matching $P\cap E$ on the vertices of the coalition. Let $M\subseteq E$ be an arbitrary matching; we say that $P$ is a (weakly/strongly) blocking path or cycle for the matching $M$ if $\mathcal{P}$ is a weakly/strongly blocking coalition for $M$, as witnessed by the matching $P\cap E$.
@@ 2.1 Weak core
As we shall see, the problems concerning the weak core are considerably simpler than those concerning the strong core, except for the optimization problem. We begin with a characterization result for the weak core.
@@
If $M$ is in the weak core, then conditions a) and b) hold, since otherwise the $M_{0}$-alternating path or cycle would block $M$. Suppose a) and b) hold, but $M$ is not in the weak core. There is a coalition of players with a matching $M^{\prime}$ such that every player in the coalition is strictly better off with $M^{\prime}$, therefore the players who have both of their vertices covered by $M$ cannot belong to the coalition – we can delete the vertices of these players from the graph, and consider only the remaining graph. Take the union of the remaining player edges and $M^{\prime}$. This consists of $M_{0}$-alternating cycles and paths. If there is an $M_{0}$-alternating cycle in the union, then it contradicts a). If there is an $M_{0}$-alternating path, then the player edges at the two ends of the path have one vertex covered by $M^{\prime}$, therefore they have 0 vertices covered by $M$, which contradicts b). ∎
@@
NTU-PMG-WC-Verif-2 is solvable in $\mathcal{O}(|V|^{4.5})$ time.
@@
We show that we can check the conditions of Lemma 2 in $\mathcal{O}(|V|^{4.5})$ time (utilizing that a maximum size matching can be computed in $\mathcal{O}(|V|^{2.5})$ time [19]). Let $M^{\prime}_{0}\subseteq M_{0}$ be the set of player edges that have at most one vertex in $V(M)$. Let $G^{\prime}$ denote the graph obtained by restricting $G$ to $V(M^{\prime}_{0})$, and adding $M^{\prime}_{0}$ to the edge set.
We can check if a given player edge $e_{A}$ is in an $M^{\prime}_{0}$-alternating cycle, since it is equivalent to checking if, after the deletion of this edge from $G^{\prime}$, the remaining graph has a perfect matching. Indeed, the remaining player edges form a matching that covers all but two vertices of $G^{\prime}$ (an almost perfect matching), and there is an augmenting $M^{\prime}_{0}$-alternating path between the two vertices of $A$ if and only if $e_{A}$ is in an $M^{\prime}_{0}$-alternating cycle in $G^{\prime}$. This shows that condition a) can be checked in $\mathcal{O}(|V|^{3.5})$ time, by repeating the above for all player edges in $M^{\prime}_{0}$.
We can also check whether there is an $M^{\prime}_{0}$-alternating path between two given player edges $e_{A}$ and $e_{B}$ that have 0 vertices covered by $M$ and that do not belong to an $M^{\prime}_{0}$-alternating cycle. To do this, we remove $e_{A}$ and $e_{B}$ from $G^{\prime}$, and check if the remaining graph has an almost perfect matching. If such a matching exists, then there is an augmenting $M^{\prime}_{0}$-alternating path between two of the four vertices of $A\cup B$, and this path cannot connect the two vertices of $A$ or the two vertices of $B$, because then $e_{A}$ or $e_{B}$ would be in an $M^{\prime}_{0}$-alternating cycle in $G^{\prime}$. Thus, condition b) is violated. Conversely, if there is no almost perfect matching after the deletion of $e_{A}$ and $e_{B}$, then there is no $M^{\prime}_{0}$-alternating path where the two ends are $e_{A}$ and $e_{B}$. Therefore, condition b) can be checked by iterating through all possible choices of $A$ and $B$ in time $\mathcal{O}(|V|^{4.5})$. ∎
The next natural problem would be to decide whether the weak core is empty. However, it turns out that the weak core of the NTU matching game with couples cannot be empty.
@@
The weak core is never empty, hence NTU-PMG-WC-Exist-2 can be solved in $\mathcal{O}(1)$ time. Furthermore, a matching in the weak core can be found in $\mathcal{O}(|V|^{3.5})$ time.
@@
We can construct a matching $M^{*}$ in the weak core the following way. Let $M_{0}$ be the perfect matching formed by the player edges as before. Iterating through the player edges, we check if there is an $M_{0}$-alternating cycle in the current graph containing that edge. If there is, then let the edges of $E$ in the cycle belong to $M$, and delete the vertices of the cycle from the graph. Repeat this with the remaining graph, until there are no $M_{0}$-alternating cycles. Note that each player edge will be checked at most once, as deleting edges cannot create an $M_{0}$-alternating cycle. Hence, this can be done in $\mathcal{O}(|V|^{3.5})$ time with maximum matching algorithms.
Let $M$ denote the resulting matching, and let $M^{*}$ be a maximum size matching in $G$ that covers every vertex of $V(M)$. This can be computed in $\mathcal{O}(|V|^{2.5})$ time. We claim that $M^{*}$ is in the weak core, since the conditions of Lemma 2 are met. Condition a) clearly holds, since the construction deleted players that were covered twice by $M$ (and hence by $M^{*}$), and the remaining graph did not contain an $M_{0}$-alternating cycle.
@@
Since the weak core is always non-empty, it is natural to consider the problem of finding a minimum-cost matching in the weak core according to some linear cost function. However, this problem turns out to be hard even for players of size one, as the following theorem shows (formally, we can add an isolated vertex to each player in the theorem and the proof below if we want to get an instance with couples).
Theorem 5.
The following problem is NP-complete even for players of size 1: given an instance $(G,\Pi)$ of NTU-PMG-WC-Exist-2, a cost vector $c\in\mathbb{Z}^{E}_{+}$, and a positive integer $k$, find a matching with cost at most $k$ in the weak core.
Proof.
Observe that if each player has size one, then a matching $M$ is in the weak core if and only if $V(M)$ is a vertex cover of the graph. This enables us to reduce Vertex Cover to our problem as follows. Let $(G=(V,E),k)$ be an instance of Vertex Cover, where we want to decide if there is a vertex cover of $G$ of size at most $k$. We construct a graph $G^{\prime}$ by adding new vertices $v_{1},\dots,v_{k}$, and adding edges from each $v_{i}$ to every vertex in $V$. The original edges have cost $k+1$, while the new edges have cost 1.
We claim that $G$ has a vertex cover of size at most $k$ if and only if the weak core of $G^{\prime}$, with players of size 1, contains a matching of cost at most $k$. Indeed, a matching of $G^{\prime}$ with cost at most $k$ cannot contain any original edges and at most $k$ new edges, and such a matching is in the weak core if and only if it contains exactly $k$ new edges and their endpoints in $V$ form a vertex cover of $G$. ∎
@@ 2.2 Strong core
A matching $M$ is in the strong core if and only if
@@
    every player edge that is in an $M_{0}$-alternating cycle has both its vertices covered by $M$,
@@
    there is no $M_{0}$-alternating path from a player edge that has none of its vertices covered by $M$ to a player edge that has at most one vertex covered by $M$,
@@
It is easy to see that if $M$ is in the strong core, then these conditions hold, since otherwise the $M_{0}$-alternating path or cycle would block $M$. Suppose that the conditions hold, but the matching $M$ is not in the strong core. Then there is a weakly blocking coalition of players, together with a matching $M^{\prime}$, such that there is one player that is better off with $M^{\prime}$ than $M$, and the others are not worse off. Take the union of $M^{\prime}$ and $M_{0}$. This is a disjoint union of $M_{0}$-alternating cycles and paths. If the player that is better off by $M^{\prime}$ has 0 vertices covered by $M$, then it has at least one vertex covered by $M^{\prime}$, so it is either in an $M_{0}$-alternating cycle, which contradicts condition a), or it is in an $M_{0}$-alternating path. In the latter case, the two player edges at the ends of the path are players who have one vertex covered by $M^{\prime}$, so they have at most one vertex covered by $M$, which contradicts condition b). If the player that is better off by $M^{\prime}$ has one vertex covered by $M$, then it has two vertices covered by $M^{\prime}$, so it is either in an $M_{0}$-alternating cycle, which contradicts condition a), or it is in an $M_{0}$-alternating path (and it is not at the end of the path), but the two player edges at the ends of the path are players that have one vertex covered by $M^{\prime}$, so they have at most one vertex covered by $M$, which contradicts condition b) or c). ∎
@@
NTU-PMG-SC-Verif-2 can be solved in $\mathcal{O}(|V|^{5.5})$ time.
@@
We need to show that we can check the conditions given by Lemma 6 in polynomial time. The techniques used here are similar to the proof of Theorem 3, so we skip some details. We can check if a given player edge is in an $M_{0}$-alternating cycle, since it is equivalent to checking if, after the deletion of this edge, the remaining graph (the original graph $G$ together with the remaining player edges) has a perfect matching. We can do this for every player edge with at most one vertex in $V(M)$.
@@
We can check whether there is an $M_{0}$-alternating path between two given player edges, knowing that these do not belong to an $M_{0}$-alternating cycle, since this is equivalent to checking if by deleting these two edges the remaining graph has an almost perfect matching. We can do this for any two player edges such that one of them has 0 vertices covered in $V(M)$, and the other one has at most one vertex in $V(M)$.
@@
For three given player edges such that none of them belong to an $M_{0}$-alternating cycle, we can check if there is an alternating path that contains all three of them: this is equivalent to checking if, by deleting all three of these edges, the remaining graph has an almost perfect matching. We can check this for any three player edges that have one vertex in $V(M)$. This leads to a runtime of $\mathcal{O}(|V|^{5.5})$. ∎
2.2.1 Deciding the existence of a strong core matching